OpenStudy (anonymous):

simplify the rational expression. state any restrictions on the variable. t^2+t-6/t^2-9

OpenStudy (anonymous):

Would you like me to walk you through it or just answer it?

OpenStudy (anonymous):

well its kinda late so it would be really helpful if you told me.

OpenStudy (anonymous):

Okie dokie, (t^2 + 2t - 24)/(t^2 - 36) = (t + 6)(t - 4)/(t + 6)(t - 6) = (t -4)/(t - 6), t is not equal to 6 Theres the quick and important part.

OpenStudy (anonymous):

sooo the answer is its not equal to 6?

OpenStudy (anonymous):

Nope, is ther more I should be adding?

OpenStudy (anonymous):

i dont think so...uh what is the exact answer?

OpenStudy (anonymous):

Can you post the answer set? I work better that way

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Okay hold on, since I can't go though each answer checking it... hows t-4/t-6?

OpenStudy (anonymous):

thats fine.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Okay thanks and have a good day ;D